OpenStudy (drummergirl7):

I NEED YOUR HELP BAD!! PLEASE HELP!!! I have 3 questions that i need answers fast! Which equations show the associative property? Select each correct answer. 3.81+(−4.81+(−6.15))=(3.81+(−4.81))+(−6.15)3.81+(−4.81+(−6.15))=(3.81+(−4.81))+(−6.15) (−7.5+(−1.2))+(−3.7)=−(7.5+1.2)+(−3.7)(−7.5+(−1.2))+(−3.7)=−(7.5+1.2)+(−3.7) (−1.5+(−8.6))+(−1.4)=−1.5+(−8.6+(−1.4))(−1.5+(−8.6))+(−1.4)=−1.5+(−8.6+(−1.4)) (−4.8+(−0.65))+1.4=−(4.8+0.65)+1.4

OpenStudy (drummergirl7):

@mathstudent55 @AloneS

OpenStudy (drummergirl7):

@Mehek14

OpenStudy (drummergirl7):

hii

OpenStudy (mathstudent55):

Do you know what the associative property is?

OpenStudy (drummergirl7):

its were u can add no matter how the numbers are and the parenthises

OpenStudy (mathstudent55):

Are you missing an equal sign in the first option?

OpenStudy (drummergirl7):

i copyed it from my school

OpenStudy (mathstudent55):

The associative tells you that when you add numbers together, the way you group the numbers together does not alter the result. This is the typical way the associative property is stated: \((a + b) + c = a + (b + c)\) That shows you that if you group the \(a\) and \(b\) together and add them, and then you add \(c\), you will get the same result as if you group together \(b\) and \(c\) and add them together, and then add \(a\) to it.

OpenStudy (drummergirl7):

i really need help. i know this is hard

OpenStudy (mathstudent55):

Sorry, the equal sign is not red, but it is big.

OpenStudy (drummergirl7):

ok

OpenStudy (mathstudent55):

Look at the first choice now corrected with the equal sign that was missing.

OpenStudy (drummergirl7):

ok now what do i do :)

OpenStudy (drummergirl7):

is it a and c? :)

OpenStudy (mathstudent55):

I'm writing each piece between equal signs in a single line to make it easier to read. The entire thing below is option A of your answers. I also numbered the lines. 3.81+(−4.81+(−6.15)) Line 1 = (3.81+(−4.81))+(−6.15) Line 2 = 3.81+(−4.81+(−6.15)) Line 3 = (3.81+(−4.81))+(−6.15) Line 4

OpenStudy (drummergirl7):

by the way thank u for all ur help u are amazing

OpenStudy (drummergirl7):

its defenitly A now we have to figure out the others

OpenStudy (eiwoh):

Aren't math questions confusing?

OpenStudy (drummergirl7):

yes

OpenStudy (mathstudent55):

Look at Line 1. -4.81 and -6.15 are in parentheses to be added together Then afterwards, you'd add 3.81 to the sum. Now look at Line 2. Now 3.81 and -4.81 are in parentheses grouped together. You'd add those first, then add -6.15 A change in grouping in an addition is what the associative property allows you to do. Option A clearly shows the associative property from Line 1 to Line 2.

OpenStudy (eiwoh):

Good job @mathstudent55 you are really good at math

OpenStudy (mathstudent55):

Now let's look at the other options to see if any other also shows the associative property being used.

OpenStudy (eiwoh):

Can I tag you when I need help on math questions? lol

OpenStudy (mathstudent55):

@Eiwoh Thanks

OpenStudy (drummergirl7):

ok so A is right now lets look at the other options

OpenStudy (mathstudent55):

In option B, a negative sign was factored out. That is the same as factoring -1 one. That is the use of the distributive property, not associative.

OpenStudy (drummergirl7):

yea i didnt think b was right

OpenStudy (drummergirl7):

i dont think D is right but i could be wrong

OpenStudy (mathstudent55):

Here is option C. \((−1.5+(−8.6))+(−1.4)=−1.5+(−8.6+(−1.4))\) Look at the left side: -1.5 and -8.6 are added in parentheses. Add them first. Then add -1.4 Now look at the right side: -8.6 and -1.4 are in parentheses. Add them first. then add -1.5 to that. Does this remind you of the associative property?

OpenStudy (mathstudent55):

You are correct about D. D is the distributive property just like B was.

OpenStudy (mathstudent55):

What about C?

OpenStudy (drummergirl7):

yes it is..... i think..

OpenStudy (mathstudent55):

@Eiwoh Yes, you can certainly tag me.

OpenStudy (mathstudent55):

Correct. C is again change of grouping just like A was. Answer: A and C.

OpenStudy (mathstudent55):

Good job @DrummerGirl7

OpenStudy (drummergirl7):

yayyyy!! hey can u help me with 2 more i promise it wont take long and i really need u

OpenStudy (drummergirl7):

@mathstudent55

OpenStudy (mathstudent55):

Ok, but pls start a new post for each question.

OpenStudy (drummergirl7):

ok